Given the following demand function: Q= 120-4P+0.01M, where M is income and stands at $400 and the prevailing market price is $10.
a) Calculate the consumer surplus in this market.
B) what is the market's willingness to pay.
1
Expert's answer
2020-09-29T09:13:19-0400

a)

"Q=120-4P+0.01M\\
But\ P=10, M=400\\
\implies 120-4(10)+0.01(400)\\
=84\\
Now\ let's\ get\ the\ choke\ price;\\
let\ Q=0\\
\therefore\ 0=84-4P\\
P=21"

Then we now have a triangle with a height of "21-10=11\ and\ base is\ 84;\\"

Consumer surplus"=\frac{1}{2}*Height\ *Base\\"

"=\frac{1}{4}*11*84\\
=462"

The height shows the willingness to pay of 11.